# Generative AI for Physical-Layer Authentication

### v2

@@ III-A2 Fingerprint Augmentation
GAI can augment fingerprint datasets, thereby enhancing the robustness of PLA models. For instance, Jiang et al. [6] employed conditional GANs to synthesize fingerprint samples for Low Earth Orbit (LEO) satellites. This conditional approach enables the synthesis of more realistic fingerprints, as environmental factors are explicitly incorporated into the generation process. Similarly, Chi et al. [12] employed GDMs to jointly leverage temporal, spectral, and complex-valued features of fingerprint waveforms, thus enhancing the generation performance of fingerprints.
@@ III-A3 Fingerprint Filling
In real-world, fingerprint data loss may occur due to physical obstructions, device malfunctions, and environmental disturbances. Such missing fingerprint data directly disrupts the spatiotemporal continuity of fingerprint sequences, necessitating fingerprint filling methods to restore complete fingerprint sequences. DAI-based methods establish data correlations only between adjacent nodes through local neighborhood modeling, making them ill-suited to adapt to dynamic changes in node positions or obstruction patterns. In contrast, GAI models, such as Transformers, leverage self-attention mechanisms to automatically learn dependency strengths between different fingerprint data points, dynamically capturing long-range spatiotemporal correlations, thereby achieving more robust missing fingerprint reconstruction performance. For example, Zhao et al. [11] introduced a Bidirectional Encoder Representations from Transformers (BERT)-based framework for robust fingerprint filling.
@@ III-A4 Fingerprint Transmission
Compared with centralized PLA, collaborative PLA utilizing multi-party cooperation exhibits superior authentication performance. As the number of dimensions of fingerprint samples increases, a key challenge is how to efficiently transmit the fingerprints estimated by collaborators to Bob. Semantic communication uses joint source-channel coding to understand the semantic content of information rather than just focusing on accurate symbol transmission, thereby reducing bandwidth requirements and potentially improving the efficiency and reliability of fingerprint transmission. GAI can help further enhance the performance of fingerprint transmission based on semantic communication, such as the construction of fingerprint knowledge bases and the management of fingerprint transmission models.
@@ III-B1 Attack Detection
GAI, leveraging self-supervised learning mechanisms, models the probability distribution of fingerprints without relying on manually annotated “fingerprint-identity” mapping relationships. Although it was not originally designed for classification tasks, through architectural adjustments, it remains capable of performing unsupervised attack detection. For example, Meng et al. [3] combined the advantages of AEs and VAEs to obtain superior attack detection performance when processing fine-grained, high-dimensional fingerprint samples. Additionally, by modifying the discriminator’s functionality to output category probability distributions while distinguishing data authenticity, unsupervised classification can be achieved using GANs. Moreover, by comparing generation probabilities under different identity category hypotheses, GDM can implicitly construct detection boundaries without explicit identity labels.
@@ III-B2 Multiuser Authentication
In communication systems, PLA models need to authenticate multiple legitimate users simultaneously and detect unknown attackers. Open-Set Recognition (OSR) provides an effective solution to this challenge. OSR requires the model to not only correctly classify known classes seen during training, but also effectively recognize and reject samples of unknown classes during the testing phase. Compared with traditional OSR methods, such as classifier-based adjustment approaches and distance metric methods, GAI has the following advantages: excelling in detecting high-dimensional unknown fingerprint samples, offering more direct detection evidence through reconstruction errors, and demonstrating stronger adaptability to unknown fingerprint categories. For instance, Guo et al. [5] introduced GANs to distinguish between legitimate and anomalous devices.
@@ III-B3 Efficient Training and Deployment
Channel Knowledge Map (CKM) helps communication systems predict CSI and reduce the overhead of real-time fingerprint estimation by constructing a location-specific CSI database. By building CKM, CSI-based downstream tasks such as PLA can be effectively empowered. Compared with traditional CKM construction methods, GAI can reduce dependence on measured CSI data, generate CSI fingerprints for specific locations, and achieve efficient transfer of different channel models. For instance, Fu et al. [4] treated incomplete CKM images as conditions for training a conditional denoiser within the decoupled GDM, thereby addressing the pivotal challenge of reconstructing a comprehensive map for all relevant locations using only partially observed channel knowledge data.
@@ III-C1 Flexibility for Eliminated Legitimate Users
As the departure of old users, GAI-based PLA models need to forget the fingerprint characteristics of old users or incorporate the relevant distributions of new users to optimize and update authentication parameters. Four machine unlearning strategies can be applied to forget the fingerprint characteristics of old users, including gradient-based optimization, knowledge distillation, data sharding, and adding learnable layers strategies.
@@ III-C2 Defense Against Intelligent Attacks
Attackers can generate adversarial samples to deviate the fingerprint features extracted at the receiver from the true distribution, thereby deceiving PLA models [13]. Additionally, attackers can embed fingerprint data with specific trigger patterns before the training phase, and activate this backdoor during the authentication phase [2]. Papangelo et al. [13] verified that adversarial training can defend against adversarial attacks for PLA models. Additionally, adversarial training can also defend against backdoor attacks using the discriminator. Moreover, VAEs can identify abnormal samples through reconstruction errors. Besides, the denoising process of GDM models can be used to eliminate fingerprint triggers, disrupting the structural consistency of the triggers.
@@ III-C3 Adaptation for Different Scenarios
GAI, especially Transformer-based LLMs, with its powerful generalization ability and multi-scenario knowledge transfer, can flexibly adapt to the requirements of PLA in different application scenarios. In the pre-training phase, LLM learns fingerprint feature distributions from various scenarios. When encountering new scenarios, LLM adopts parameter-efficient fine-tuning techniques to adapt to local fingerprint features by updating some parameters, thereby avoiding the computational cost of retraining the entire model. For instance, Liu et al. [9] combined LLMs and proposed a BERT-based lightweight PLA framework to enhance the zero-trust edge IoT security.
IV Case Study: GAI-Based PLA for Dynamic Environments
Figure 2: Illustration of the proposed GDM-based PLA scheme for dynamic environments. During offline training, Alice’s fingerprints $X_{A}$ and Jack’s fingerprints $X_{J}$ are jointly pre-processed as one image $X$, and GDM is utilized to learn the joint fingerprint distributions. During online authentication, the trained GDM is leveraged to output predicted Alice’s fingerprints $\hat{X}_{A}$ according to Jack’s fingerprints, and the identities of estimated unknown fingerprints $\tilde{X}$ are obtained by comparing them with $\hat{X}_{A}$.
IV-A Proposed GDM-based PLA Scheme
IV-A1 Challenges
Existing PLA schemes based on CSI fingerprints face challenges in dynamic wireless environments, where sudden environmental fluctuations cause drastic changes in the spatial distribution and temporal dependence of CSI fingerprints. These rapid changes make the previously learned decision boundaries no longer applicable to the current channel environment, thereby constraining authentication performance.
IV-A2 Motivations
@@
Based on channel extrapolation principles [14], we hypothesize correlations between the CSI fingerprints of the legitimate transmitter (Alice) and collaborator (Jack). We further employ GDM to learn the association between their CSI fingerprints. During authentication, by leveraging Jack’s fingerprints as the condition, the trained GDM can improve the prediction accuracy of Alice’s fingerprints, thus enhancing the authentication performance.
IV-A3 System Model
As shown in Figure 2, Alice is a legitimate device that requires identity authentication performed at Bob, Jack is a collaborative device, and Eve is a spoofing attacker impersonating Alice. We assume that Alice and Jack are both stationary, while the surrounding environment changes rapidly. The fixed and close distance between Alice and Jack allows for the establishment of a stable relationship between their CSI fingerprints through channel extrapolation in the user domain [14].
IV-A4 Steps of the Proposed Scheme
As illustrated in Figure 2, the proposed scheme includes the following steps:
  • Offline Training: Bob obtains Alice’s fingerprints $X_{A}$ and Jack’s fingerprints $X_{J}$ through channel estimation. Their distributions are further learned through GDM. The problem is formulated as fitting a function $X_{A}=f(X_{J})$ to predict $X_{A}$ using $X_{J}$. Here, we train GDM to represent $f$. Specifically, it gradually adds noise to disrupt fingerprint data, and then learns to recover the original data from the noise through a reverse denoising process. The loss function of the denoising process is $\mathcal{L}=\mathbb{E}_{X_{A},X_{J},\epsilon_{t},t}\left\|\epsilon_{t}\;-\;\epsilon_{\Theta}\bigl{(}X_{A,t},\,t,\,X_{J}\bigr{)}\right\|_{2}^{2}$, where $\epsilon_{t}$ is the Gaussian noise added to $X_{A}$ at diffusion step $t$, $X_{A,t}$ is the noisy fingerprint after $t$ diffusion steps, and $\epsilon_{\Theta}(X_{A,t},t,X_{J})$ is the predicted noise by model parameters $\Theta$.
@@ IV-A4 Steps of the Proposed Scheme
  • Online Authentication: Bob obtains Jacks’ fingerprints $X_{J}$, and leverages the trained GDM to predict Alice’s fingerprint $\hat{X}_{A}$. Then, Bob estimates fingerprints $\tilde{X}$ from unknown transmitters, and predict the identities by comparing the differences between $\hat{X}_{A}$ and $\tilde{X}$.
@@
Figure 3: Simulation parameters and simulation results. During training, we have 10, 860 fingerprint samples for Alice and Jack. During authentication, we run 50 times to obtain the average result, and we have 50 fingerprints for Alice, Eve, and Jack in each time. We use $F_{1}=(2P_{tl}P_{ta})/(P_{tl}+P_{ta})$ and authentication error rate $R_{e}$ to evaluate the authentication performance [3]. $F_{1}\in[0,1]$, and larger $F_{1}$ means better authentication performance. $P_{tl}$ and $P_{ta}$ respectively represent correct authentication rate and false alarm rate. $R_{e}\in[0,1]$, and larger $R_{e}$ means worse authentication performance. The proposed scheme utilizes GDMs to generate Alice’s fingerprint at time $T$ based on Jack’s fingerprint at time $T$, and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 1: The fingerprint to be authenticated at time $T$ is compared with Alice’s fingerprint at time T1T-1. Baseline Scheme 2: The typical time series model, Long Short-Term Memory (LSTM) model, is used to predict Alice’s fingerprint at time $T$ and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 3: The fingerprint to be authenticated at time $T$ is compared with Jack’s fingerprint at time $T$.
IV-A5 Complexity Analysis
@@
The computational complexity is denoted as $\mathcal{O}\left(TBLN\left(2C^{2}{S}+4{S}^{2}C\right)\right)$, where $T$ is the total time step, $B$ is the batch size, $L$ is the number of layers, $N$ is the number of ResNetBlocks in each layer, and $C$ and S{S} denote the maximum channel number and spatial resolution, respectively. $2$ reflects the dual-branch design, while $4$ in the attention term arises from the addition of cross-attention alongside self-attention in every block. According to the simulation results on the computer configured as an NVIDIA RTX A6000 GPU paired with an Intel Xeon w7‑2495X CPU, each time step requires 18.05 ms.
IV-B Simulation Results
The DeepMIMO datasets are employed to generate CSI fingerprint samples for simulation experiments. To preserve the phase integrity of complex fingerprints, we compute the dynamic ranges of its real and imaginary components as the differences between their maxima and minima, and then use the larger of these two ranges as a common normalization factor to linearly map both the real and imaginary values into the interval [1,1][-1,1]. Figure 3 illustrates the DeepMIMO parameters, GDM parameters, the training process, and the simulation results. The proposed scheme can accurately recognize the identity of each fingerprint during authentication, but the three baseline schemes can not. Due to the time-varying nature of wireless channels and the continuous change in Alice’s position, there are deviations in the distribution of Alice’s channel fingerprints between the training and authentication phases. The proposed scheme can learn the relative features between Alice’s and Jack’s fingerprints through GAI and predict Alice’s fingerprints based on Jack’s fingerprints during authentication, thereby improving authentication performance in dynamic environments. In contrast, baseline scheme 1 only utilizes Alice’s fingerprints from the previous moment, which differs from the current distribution of Alice’s fingerprints. Although baseline scheme 2 employs LSTM to learn the distribution changes of fingerprints, its ability to learn fingerprint distributions is weaker than that of GAI. Because there are correlations between Alice’s and Jack’s fingerprint distributions, baseline scheme 3 can obtain better authentication performance in low SNR environments than baseline schemes 1 and 2.
IV-C Lessons Learned
The case study leverages channel extrapolation principles and GDMs to model and learn the joint distributions of Alice’s and collaborator’s channel fingerprints. Due to the excellent generation ability and noise robustness of GDMs, the authentication performance of this scheme in dynamic environment is significantly better than that of the three baseline schemes.
V Challenges and Future Directions
V-A Challenges
V-A1 Computational Demands and Latency Implications
Although GAI can enhance the performance of PLA, its high computing power requirements also bring some challenges to its deployment. Currently, the growth of GAI model sizes has outpaced advancements in general-purpose compute hardware. Both parameter counts and per-request floating-point operations (FLOPs) are scaling more rapidly than accelerator performance improvements, creating escalating compute footprints for both training and inference. Additionally, the inference time of GAI-based PLA models will have a significant impact on the communication latency. Especially in the dynamic environment, with the continuous change of fingerprint identity characteristics, high-frequency identity authentication is required to ensure the legitimacy of identity.
V-A2 Security of GAI-based PLA Models
During the training process of GAI model, it directly contacts and remembers the features of fingerprint data. Intelligent attackers may launch member inference attacks to determine whether a specific fingerprint data exists in the training dataset. They could further reverse engineer privacy sensitive fingerprint data such as the location of legitimate users. Additionally, attackers may tamper with the training fingerprint data of GAI models through adding adversarial data, causing the GAI model to learn incorrect fingerprint distributions. Attackers may also implant backdoors during training, which hide a mapping relationship between triggers and malicious outputs. When encountering specific triggering patterns, these backdoors may trigger pre-defined error outputs.
V-A3 Trade-off Issues
The trade-off between security performance and complexity in GAI-based PLA is essentially about improving security through GAI technology while controlling the computational, data, and hardware overhead introduced as a result. As mentioned above, GAI can enhance the fingerprint collection, model training, and performance optimization stages of PLA. Although GAI technology improves access security, it brings challenges such as computational complexity and hardware requirements, which increase deployment costs. In addition, GAI may introduce new security threats, such as adversarial attacks, which also require sacrificing complexity to further enhance the reliability of GAI models.
V-A4 GAI-based PLA for Emerging Scenarios
As a typical scenario for 6G, ISAC achieves the goal of spectrum conservation and mutual enhancement between communication and sensing. However, there are inconsistencies in the spatiotemporal dimensions of channel fingerprints, RF sensor fingerprints, and non-RF sensor data. Additionally, to support secure access for massive communication in 6G, PLA models need to pre-register and learn the fingerprint features of a massive number of devices. Besides, 6G will also support Hyper Reliable and Low-Latency Communication (HRLLC). However, GAI models typically require high computational resources. Thus, the inference time may exceed the latency budget, especially in resource constrained IoT devices, which cannot meet the requirements of real-time PLA.
V-B Future Directions
V-B1 Model Compression for GAI-based PLA
Model compression can not only reduce the complexity of GAI models, but also potentially defend against new security threats such as backdoor attacks. Model compression techniques mainly include model pruning, parameter quantization, low-rank decomposition, and knowledge distillation. For example, knowledge distillation transfers the knowledge of the teacher model to the student model while preserving core features, and model pruning can reduce erroneous outputs in specific input fingerprint patterns by removing redundant neurons or connections, thus removing implanted backdoor triggering features.
V-B2 Federated Learning for GAI-based PLA
Federated learning [15] provides a foundational framework for trusted training of GAI-based PLA models through a distributed training architecture. It allows legitimate users to collaborate in training global GAI models without sharing locally collected fingerprint datasets. It can be combined with differential privacy to further enhance the privacy protection capability of GAI-based PLA models.
V-B3 Distributed Trust for GAI-based PLA
Distributed trust builds a decentralized trust mechanism to ensure the credibility of the behavior of each distributed node and the overall security of the system. By using distributed ledger technology to record the source of training fingerprint data for GAI, it ensures that the fingerprint data is traceable and tamper proof. Additionally, multi-party secure computing (MPC) or trusted execution environment (TEE) can be combined to process fingerprint samples in a distributed environment, protecting user privacy.
V-B4 Emerging GAI Models for PLA
Some emerging GAI models, such as Large Language Models (LLMs), can be introduced to further enhance the performance of PLA. By utilizing the powerful representation and generation potential of LLMs, an efficient and dynamically evolving fingerprint database can be constructed. LLMs can generate fingerprints in different physical environments, covering extreme scenarios and solving the problem of scarce fingerprint data. In addition, by inputting environmental parameters, LLM can generate fingerprints that are suitable for the current environment. Moreover, LLMs can be used to generate semantic labels for fingerprints to support fast retrieval and matching of fingerprints.
V-B5 GAI-based PLA for 6G
GAI plays a critical role in learning multi-modal fingerprints for ISAC. It constructs a unified cross-modal representation space through latent space mapping, effectively resolving the alignment challenges posed by heterogeneous data. Additionally, in scenarios where partial sensors fail, the cross-modal reasoning capabilities of GAI models enable the reconstruction of consistent representations for missing modalities, ensuring the completeness of the authentication system’s input. Additionally, LLMs are promising to support secure access in massive communication and ubiquitous connectivity. By fine-tuning the pre-trained LLMs, the customized authentication needs of specific scenarios can be met.

### v3

@@ III-A2 Fingerprint Augmentation
GAI can augment fingerprint datasets, thereby enhancing the robustness of PLA models. For instance, Jiang et al. [5] employed conditional GANs to synthesize fingerprint samples for Low Earth Orbit (LEO) satellites. This conditional approach enables the synthesis of more realistic fingerprints, as environmental factors are explicitly incorporated into the generation process.
@@ III-A3 Fingerprint Filling
In the real-world, fingerprint data loss may occur due to physical obstructions, device malfunctions, and environmental disturbances. Such missing fingerprint data directly disrupts the spatiotemporal continuity of fingerprint sequences, necessitating fingerprint filling methods to restore complete fingerprint sequences. DAI-based methods establish data correlations only between adjacent nodes through local neighborhood modeling, making them ill-suited to adapt to dynamic changes in node positions or obstruction patterns. In contrast, GAI models, such as Transformers, leverage self-attention mechanisms to automatically learn dependency strengths between different fingerprint data points, dynamically capturing long-range spatiotemporal correlations, thereby achieving more robust missing fingerprint reconstruction performance [15].
@@ III-A4 Fingerprint Transmission
Collaborative PLA utilizes multi-party cooperation to exhibit superior authentication performance. As the dimensionality of fingerprint samples increases, a key challenge is how to efficiently transmit the fingerprints estimated by collaborators to Bob. Semantic communication uses joint source-channel coding to understand the semantic content of information rather than just focusing on accurate symbol transmission, thereby reducing bandwidth requirements. It can potentially improve the efficiency and reliability of fingerprint transmission. GAI can further help enhance the performance of fingerprint transmission based on semantic communication, such as the construction of fingerprint knowledge bases and the management of fingerprint transmission models.
@@ III-B1 Attack Detection
GAI, leveraging self-supervised learning mechanisms, models the probability distribution of fingerprints without relying on manually annotated “fingerprint-identity” mapping relationships. Although it was not originally designed for classification tasks, through architectural adjustments, it remains capable of performing unsupervised attack detection. For example, Meng et al. [9] combined the advantages of AEs and VAEs to obtain superior attack detection performance when processing fine-grained, high-dimensional fingerprint samples.
@@ III-B2 Multiuser Authentication
Open-Set Recognition (OSR) provides an effective solution for authenticating multiple legitimate users simultaneously and detecting unknown attackers. OSR requires the model to not only correctly classify known classes seen during training, but also effectively recognize and reject samples of unknown classes during the testing phase. Compared with traditional OSR methods, such as classifier-based adjustment approaches and distance metric methods, GAI has the following advantages: excelling in detecting high-dimensional unknown fingerprint samples, offering more direct detection evidence through reconstruction errors, and demonstrating stronger adaptability to unknown fingerprint categories [3].
@@ III-B3 Efficient Training and Deployment
Channel Knowledge Map (CKM) helps predict CSI fingerprints and reduce the overhead of real-time fingerprint estimation by constructing a location-specific CSI database. Compared with traditional CKM construction methods, GAI can reduce dependence on measured CSI data, generate CSI fingerprints for specific locations, and achieve efficient transfer of different channel models. For instance, Fu et al. [2] treated incomplete CKM images as conditions for training a conditional denoiser within the decoupled GDM, thereby addressing the pivotal challenge of reconstructing a comprehensive map for all relevant locations using only partially observed channel knowledge data.
@@ III-C1 Flexibility for Eliminated Legitimate Users
With the departure of old users, GAI-based PLA models need to forget the fingerprint characteristics of old users or incorporate the relevant distributions of new users to optimize and update authentication parameters. Four machine unlearning strategies can be applied to forget the fingerprint characteristics of old users, including gradient-based optimization, knowledge distillation, data sharding, and adding learnable layers strategies.
@@ III-C2 Defense Against Intelligent Attacks
Attackers can generate adversarial samples to deviate the fingerprint features extracted at the receiver from the true distribution, thereby deceiving PLA models. Papangelo et al. [10] established an experimental platform using Software-Defined Radio (SDRs) hardware (USRP X310) and demonstrated that over 99% of adversarial attacks can be effectively mitigated through GAN-based adversarial training.
@@ III-C3 Adaptation for Different Scenarios
GAI, especially Transformer-based LLM, with its powerful generalization ability and multi-scenario knowledge transfer, can flexibly adapt to the requirements of PLA in different application scenarios. In the pre-training phase, LLM learns fingerprint feature distributions from various scenarios. When encountering new scenarios, LLM adopts parameter-efficient fine-tuning techniques to adapt to local fingerprint features by updating some parameters, thereby avoiding the computational cost of retraining the entire model [7].
III-D Lessons Learned
As presented in Table II, we compare how representative GAI models address PLA issues from each of fingerprint collection, model training, and performance optimization stages.
  • During the fingerprint collection stage, GAI empowers fingerprint estimation, fingerprint augmentation, fingerprint filling, and fingerprint transmission to improve accuracy, enrich diversity, ensure completeness, and enhance transmission efficiency, respectively, thereby providing reliable fingerprint samples for training PLA models [12, 5, 1, 15].
  • During the model training stage, GAI-driven training strategies are tailored to specific task requirements, such as unsupervised attack detection, multi-user authentication, and efficient training and deployment, to develop PLA models [9, 3, 2].
  • During the performance optimization stage, the trained PLA model undergoes further enhancement through GAI to achieve advanced capabilities, including improving flexibility for eliminated legitimate users, enhancing robust defense against intelligent attacks, and boosting adaptability across diverse scenarios [10, 7].
IV Proposed GAI-enhanced PLA Framework for Dynamic Environments
IV-A Challenges
Existing CSI fingerprint-based PLA schemes face the following challenges in dynamic environments.
  • Rapidly Changing Fingerprint Distributions: Frequent device movement or rapid environmental changes significantly alter the distribution characteristics of CSI fingerprints, undermining the generalization of decision boundaries learned by DAI models and leading to degraded authentication performance.
  • Reliance on A Large Number of High-quality Fingerprints: Most schemes exhibit heightened sensitivity to training fingerprint quantity, scale, and spatial distribution uniformity, severely restricting the generalization capability in dynamic scenarios.
  • Channel Estimation Errors: While time-series-based models aim to address channel variations in dynamic environments, the high-dimensional and complex structural nature of CSI fingerprints renders them highly susceptible to channel estimation errors, particularly under low SNR conditions.
IV-B Framework Design
@@
Based on channel extrapolation principles [13], we hypothesize correlations between the CSI fingerprints of the legitimate transmitter and collaborator, which can be learned through GAI. By leveraging Jack’s fingerprints as the condition, the trained GAI model can improve the prediction accuracy of Alice’s fingerprints, thus enhancing the authentication performance. As illustrated in Figure 2, the proposed framework includes the following steps:
  1. 1.
    Offline Training: As illustrated in Figure 2 Part (A), it includes:
    1. (a)
      Channel Fingerprint Pre-processing: Bob obtains Alice’s fingerprints $X_{A}$ and Jack’s fingerprints $X_{J}$ through channel estimation. Then the fingerprints are pre-processed into one image representation for further joint analysis.
    2. (b)
      Joint Distribution Learning: The joint distributions are learned through GAI model training. The problem is formulated as fitting a function $X_{A}=f(X_{J})$ to predict $X_{A}$ using $X_{J}$.
  2. 2.
    Online Authentication: As illustrated in Figure 2 Part (B), it includes:
    1. (a)
      Dynamic Fingerprint Prediction: Bob obtains Jack’s fingerprints $X_{J}$, and leverages the trained GAI model to predict Alice’s fingerprint $\hat{X}_{A}$.
    2. (b)
@@
      Identity Authentication: Bob estimates fingerprints $\tilde{X}$ from unknown transmitters, and predicts the identities by comparing the differences between $\hat{X}_{A}$ and $\tilde{X}$.
IV-C Case Study: PLA based on GDM and Channel Extrapolation
Figure 2: Illustration of the proposed GAI-enhanced PLA framework for dynamic environments, where Alice is a legitimate device that requires identity authentication performed at Bob, Jack is a collaborative device, and Eve is a spoofing attacker impersonating Alice. Part (A) denotes the offline training step including channel fingerprint pre-processing and joint distribution learning, and Part (B) denotes the online authentication step including dynamic fingerprint prediction and identity authentication.
Figure 3: Illustration of GDMs used for CSI fingerprint generation and prediction. To preserve the phase integrity of complex CSI fingerprints, we compute the dynamic ranges of the real and imaginary components as the differences between their maxima and minima, and then use the larger of these two ranges as a common normalization factor to linearly map both the real and imaginary values into the interval [1,1][-1,1]. Part (A) denotes the training process of the GDM-based fingerprint generation, including the forward and denoising processes, and Part (B) denotes the visual denoising process for DeepMIMO “O1” outdoor and “I3” indoor CSI fingerprint datasets.
@@ IV-C Case Study: PLA based on GDM and Channel Extrapolation
Figure 4: Comparison results between the proposed GDM-based PLA scheme and four baseline schemes. The proposed scheme utilizes GDMs with a cross-attention mechanism to generate Alice’s fingerprint at time $T$ based on Jack’s fingerprint at time $T$, and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 1: The VAE model with a cross-attention mechanism, is used to predict Alice’s fingerprint at time $T$ and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 2: The typical time series model, Long Short-Term Memory (LSTM) model, is used to predict Alice’s fingerprint at time $T$ and compare it with the fingerprint to be authenticated at time $T$. Baseline Scheme 3: The fingerprint to be authenticated at time $T$ is compared with Jack’s fingerprint at time $T$. Baseline Scheme 4: The Gated Recurrent Unit (GRU) model, is used to predict Alice’s fingerprint at time $T$ and compare it with the fingerprint to be authenticated at time $T$. Part (A) uses $F_{1}=(2P_{tl}P_{ta})/(P_{tl}+P_{ta})\in[0,1]$ to evaluate the authentication performance, where $P_{tl}$ and $P_{ta}$ respectively represent correct authentication rate and false alarm rate [9]. A larger $F_{1}$ means better authentication performance. Part (B) employs authentication error rate $R_{e}\in[0,1]$ as the metric, and a smaller $R_{e}$ means better authentication performance.
IV-C1 System Model
As shown in Figure 2, we assume that Alice and Jack are both stationary, while the surrounding environment changes rapidly. The fixed and close distance between Alice and Jack allows for the establishment of a stable relationship between their CSI fingerprints through channel extrapolation in the user domain [13].
IV-C2 GDM Training
Here, we train the GDM to represent $f$. As illustrated in Figure 3 Part (A), it gradually adds noise to disrupt fingerprint data, and then learns to recover the original data from the noise through a reverse denoising process. The loss function of the denoising process is $\mathcal{L}=\mathbb{E}_{X_{A},X_{J},\epsilon_{t},t}\left\|\epsilon_{t}\;-\;\epsilon_{\Theta}\bigl(X_{A,t},\,t,\,X_{J}\bigr)\right\|_{2}^{2}$, where $\epsilon_{t}$ is the Gaussian noise added to $X_{A}$ at diffusion step $t$, $X_{A,t}$ is the noisy fingerprint after $t$ diffusion steps, and $\epsilon_{\Theta}(X_{A,t},t,X_{J})$ is the predicted noise by model parameters $\Theta$.
IV-C3 Complexity Analysis
@@
The computational complexity is denoted as $\mathcal{O}\left(TBLN\left(2C^{2}{S}+4{S}^{2}C\right)\right)$, where $T$ is the total time step, $B$ is the batch size, $L$ is the number of layers, $N$ is the number of ResNetBlocks in each layer, and $C$ and S{S} denote the maximum channel number and spatial resolution, respectively. $2$ reflects the dual-branch design, while $4$ in the attention term arises from the addition of cross-attention alongside self-attention in every block.
IV-C4 Simulation Results
The DeepMIMO datasets are employed to generate CSI fingerprint samples. The simulations are conducted on the computer configured as an NVIDIA RTX A6000 GPU paired with an Intel Xeon w7‑2495X CPU.
  • Inference Latency: Using the Denoising Diffusion Implicit Model (DDIM) accelerator with 20 sampling steps, the total inference time is reduced to approximately 180.6 ms.
  • Energy Consumption: Based on hardware specifications, the system exhibits a total peak power of approximately 525 W. Consequently, the estimated energy consumption for a single authentication instance (180.6 ms) is calculated at approximately 94.8 J.
  • Visualization of CSI Fingerprint Generation: As illustrated in Figure 3 Part (B), the frequency-domain CSI fingerprint generation in DeepMIMO “O1” outdoor and “I3” indoor scenarios is intuitively visualized. The comparison between the true and generated CSI fingerprints reveals that both real and imaginary fingerprint components align closely in structural distribution and amplitude range. This high-fidelity reconstruction across scenarios strongly validates GDM’s accuracy and robustness in handling complex CSI fingerprint characteristics.
  • Comparison Results: As illustrated in Figure 4, the authentication performance across varying SNRs from 5 dB to 20 dB demonstrates that, the proposed scheme consistently maintains an F1 score of 1.0 and an authentication error rate of 0. This confirms its ability to accurately predict and generate CSI fingerprints for PLA. In contrast, baseline scheme 1 is constrained by limited feature fusion capabilities, while baseline schemes 2 and 4, relying solely on time-series modeling without spatial inference, fail to effectively capture complex channel mappings, resulting in significantly inferior performance. Furthermore, baseline scheme 3 fails almost completely due to the lack of an inference mechanism, validating the necessity of combining channel extrapolation with high-fidelity generative modeling to achieve highly reliable PLA.
V Future Research Roadmap
V-A Enhanced GAI Architectures for Highly reliable
Fingerprints Existing work improves RF fingerprint generation by learning the time, frequency, and complex-valued domains [1]. To generate highly reliable fingerprints, future directions include: (1) integrating physical laws of fingerprints to ensure accuracy; (2) designing multimodal GAI models supporting cross-modal fingerprint learning in ISAC systems; and (3) combining multiple GAI models for fine-grained fingerprint processing.
V-B Lightweight GAI Design for Efficient PLA
Current approaches use knowledge distillation to reduce model size and computational complexity [8]. To enhance inference efficiency, future work should: (1) optimize GAI frameworks through efficient sampling techniques tailored for GDMs; (2) integrate quantization to implement low-precision parameters, reducing storage and computational demands; and (3) develop plug-and-play hierarchical modules enabling adaptive component selection during deployment based on resource constraints.
V-C Generalization Enhancement for GAI-based PLA
Existing methods employ data augmentation to increase fingerprint [1] or combine pre-training and fine-tuning to improve the generalization of LLM-based PLA [7]. Future research should: (1) design conditional GAI models for environment-specific fingerprint generation; (2) introduce meta-learning or incremental learning to adapt GAI models to environmental changes; (3) adopt domain adaptation to learn fingerprint distribution gaps between training and target environments; and (4) integrate digital twins to optimize the generalization of GAI models in unseen environments.
V-D Securing GAI-based PLA
Adversarial training mitigates adversarial attacks [10], but some intelligent attacks such as backdoor and poisoning attacks oriented to GAI models remain unresolved. Future directions include: (1) deploying adversarial cleaning, model pruning, robust optimization, and trusted execution environments (TEEs) to defend against backdoor and poisoning attacks across pre-processing, training, and authentication stages; (2) using machine unlearning to erase specific knowledge and counter adversarial attacks; (3) applying differential privacy, knowledge distillation, and regularization to address membership inference attacks and protect the privacy of training fingerprint samples; and (4) implementing federated learning or secure multi-party computation for distributed training [4], avoiding centralized PLA limitations.
V-E Promoting the Industrialization of GAI-based PLA
GAI-based PLA solutions for IoT [11], IIoT [9], and 6G edge intelligence [7] are developed. Industrialization requires: (1) establishing benchmark fingerprint datasets and GAI models for representative scenarios; (2) designing interdisciplinary PLA methods to leverage GAI technologies; and (3) establishing diversified evaluation standards to comprehensively assess the real-world deployment effectiveness of GAI-based PLA.